OpenStudy (anonymous):

How do I find the equation of an exponential equation with 2 points? (1,8) and (1,2) Thanks!

OpenStudy (tkhunny):

You need an exponential equation with two parameters. However, as these have the same x-value, this could be a challenge. How about \(x = e^{0\cdot y}\) - I'm going to guess that is not what is wanted, but that makes it no less correct.

OpenStudy (anonymous):

I'm not sure what you mean tkhunny? I remember the equation being y = a*B^x I believe

zepdrix (zepdrix):

Tachi, did you write down the coordinate pairs correctly? Because what you have written can not represent y as a function of x.

OpenStudy (anonymous):

Oh, wait sorry. (-1,8) and (1,2) My mistake!

zepdrix (zepdrix):

Ah ok :)

zepdrix (zepdrix):

Ok so the form that you provided looks good. We'll use the coordinate pairs to figure out our a and b.

zepdrix (zepdrix):

If we plug in the coordinate pairs, we get a system of equations.\[\Large 8=aB^{-1}\]\[\Large 2=aB^1\] Understand how I set those up?

zepdrix (zepdrix):

Crap I need to leave for class :( Lemme just rush through and give you some notes so you can solve this.

zepdrix (zepdrix):

Divide the first equation by the second.\[\Large \frac{8}{2}=\frac{aB^{-1}}{aB^1}\]The a's will cancel out, allowing you to solve for B.

zepdrix (zepdrix):

After you have found your B, plug it into this form:\[\Large y=aB^x\]Then plug in one of your coordinate pairs in order to solve for a.

OpenStudy (anonymous):

ohh, I see. Thanks so much! Those steps are extremely clear and I think I got it!